OpenStudy (anonymous):

Show the procedure for finding the transform of f(t)=cos5t-sin5t

OpenStudy (anonymous):

Fourier?

OpenStudy (anonymous):

It doesn't say but thats what I'm guessing since thats what we've covered.

OpenStudy (anonymous):

You haven't covered Laplace?

OpenStudy (anonymous):

Series and Sequences are my absolute weakness.

OpenStudy (anonymous):

I have covered LaPlace as well.

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

That was our last week of my class

OpenStudy (anonymous):

Its a short answer type question though so I'm not sure what it means by transform.. the only 2 I can recall are fourier and laplace

OpenStudy (anonymous):

When's it due?

OpenStudy (anonymous):

3 hours

OpenStudy (anonymous):

ouch

OpenStudy (anonymous):

Yeah I have like 20 more questions, but they are much more simple.

OpenStudy (anonymous):

The answer's going to depend on whether it's Fourier or Laplace, that's the problem.

OpenStudy (anonymous):

Well wouldn't fourier be a series? Which would give me an interval?

OpenStudy (anonymous):

I have to leave for a bit - bad timing. I'll try and be back before your three hours is up.

OpenStudy (anonymous):

Cool thanks dude. No problem

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

I can do it quickly if it's Laplace transform.

OpenStudy (anonymous):

I haven't done any Fourier transform in a while.

OpenStudy (anonymous):

I think its LaPlace

OpenStudy (anonymous):

I don't think we did Fourier transforms, we did Fourier Series.

OpenStudy (anonymous):

They are different right?

OpenStudy (anonymous):

yes, they are.

OpenStudy (anonymous):

Ok well I guess its a laplace because I haven't done any transform other than LaPlace

OpenStudy (anonymous):

It would be easy then. do you use a Laplace transform table? or should I solve using the definition of Laplace transform?

OpenStudy (anonymous):

We use transform tables usually....

OpenStudy (anonymous):

that's good.. we will just use two formulas, which are: \[1) L[\sin kt]={k \over s^2+k^2}\] \[2) L[\cos kt]= {s \over s^2+k^2}\] where L[f(t)] denotes the Laplace transform of f(t).

OpenStudy (anonymous):

do you think you can apply these two formulas?

OpenStudy (anonymous):

Hmm Let me see... I'm really not fluent in using the table as it was our last topic discussed in the class.

OpenStudy (anonymous):

Wouldn't I just substitute K for the cosine and sine values?

OpenStudy (anonymous):

yeah k in this case is what?

OpenStudy (anonymous):

5

OpenStudy (anonymous):

OK.. show me what you get.

OpenStudy (anonymous):

\[5-s/s^2+25\]

OpenStudy (anonymous):

but I can factor the bottom if I wanted

OpenStudy (anonymous):

are you sure about factorizing the bottom?

OpenStudy (anonymous):

Nevermind I can;t

OpenStudy (anonymous):

f(t)=5-s/s^2+25 final solution.

OpenStudy (anonymous):

1) well what is the Laplace transform of f(t)? 2) is it 5-s or s-5?

OpenStudy (anonymous):

s-5 I read it inccorectly

OpenStudy (anonymous):

ok.. the notation that's usually used to denote a function after being transformed is uppercase letter, in our case it will be: \[F(s)= {s-5 \over s^2+25}\]

OpenStudy (anonymous):

Oh right... because you're transforming it in terms of f(t) and s... I recall that now.

OpenStudy (anonymous):

if you're asking about how you can get the formulas (which is the most important part of the answer), you can just use the definition and integrate for the first one: \[\int\limits_{0}^{\infty}{\cos (kt)} e ^{-st} dt\] the result of this integral (can be integrated using integration by parts) is, by definition, the Laplace transform of cos (kt).

OpenStudy (anonymous):

Right which will help me build a transform table if I didn't have one available?

OpenStudy (anonymous):

Exactly.

OpenStudy (anonymous):

Sweet tthanks a lot :)

OpenStudy (anonymous):

You're welcome!